A triangle has a perimeter of 53 inches. The medium side is 5 inches more than the shortest side, and the longest side is twice the shortest side. Find the shortest side. Thank You
Answer by rfer(16322)

x+x+5+2x=53
4x=53-5
x=48/4
x=12
x+5=17
2x=24